### v3

# Technological Complexity Based on Japanese Patent Data

Abstract
As international competition intensifies in technologies, nations need to identify key technologies to foster innovation. However, the identification is challenging due to the independent and inherently complex nature of technologies. Traditionally, regional analyses of technological portfolios have been limited to binary evaluations, indicating merely whether a region specializes in a technology, or relying on the average Technological Complexity Index (TCI) of the specialized technologies. This study proposes that evaluating TCI at the corporate level could provide finer granularity and more detailed insights. To address the underutilization of corporate-level TCI assessments in Japan, this study applies the Technological Complexity Index using carefully processed patent data spanning fiscal years 1981 to 2010. Specifically, we analyze a bipartite network composed of 1,938 corporations connected to technological fields categorized into either 35 or 124 classifications. Our findings quantitatively characterize the ubiquity and sophistication of each technological field, reveal detailed technological trends reflecting broader societal contexts, and demonstrate methodological stability even when employing finer technological classifications. Additionally, our corporate-level approach allows consistent comparisons across different regions and technological fields, clarifying regional advantages in specific technologies. This refined analytical framework offers policymakers and researchers robust, targeted insights, thereby significantly contributing to innovation strategy formulation in Japan.
keywords:
Keyword1, Keyword2, Keyword3
Introduction
In the twenty-first century, intensifying international competition has made technological capability a central element of national competitiveness. Nations seek to foster environments that support the development of sophisticated technologies, as these capabilities are fundamental to sustainable economic growth. [1, 2, 3] The ecosystems that produce such capabilities are, however, inherently complex, involving interactions among countries, corporations, and technologies [4]. This complexity makes it difficult to accurately identify the key technologies and actors within the system. Understanding this structure is crucial for developing effective science and technology strategies to strengthen a nation’s competitive advantage.
Historically, the assessment of technological capabilities has predominantly relied on quantitative metrics derived from bibliographic data of patents and publications, ranging from simple patent counts to more complex citation-based indicators, including Technology Relevance (TR) [5, 6]. While such indicators offer certain insights by considering citation practices, patent age, and technological fields, they suffer from significant shortcomings. Their inherent dependence on the long-term accumulation of citations makes them susceptible to biases from citation behaviors, patent examiners, and strategic patenting. Crucially, these metrics fail to capture the core aspect of technological complexity, namely the intricate interplay among corporations and technologies. Consequently, approaches that rely solely on these data provide a limited perspective and inadequately reflect the qualitative dimensions of technological capability.
The HH algorithm provides a powerful solution by considering both the diversity of a region and the ubiquity of its products or technologies [7]. This dual perspective allows the methodology to distinguish between components and quantify their complexity, revealing detailed economic structures not captured by traditional metrics. The utility of this foundational framework has led to its extension into several distinct streams of research[8, 9, 10, 11, 12]. One major development is the introduction of complementary metrics such as fitness-based approaches [13, 14, 15]. Another significant stream, which is the focus of this study, involves the adaptation of the algorithm to patent data, pioneered by Balland and Rigby to measure technological complexity [16]. Building on this work, recent studies have further extended the patent-based analysis to contexts such as Asian economies [17, 18]. Despite this wide range of applications, the focus has predominantly remained at a macro level, such as on cities or regions [8]. Consequently, even within the stream of patent analysis, comprehensive validation at the corporation level is still limited.
To address this gap, our study applies the HH algorithm to an extensive set of Japanese patent records to quantify technological complexity at the corporation level. By constructing bipartite networks linking corporations to technology classes, we develop a detailed measure of technological sophistication. Our analysis investigates the relationships between technological complexity, ubiquity, and the average corporate diversity of patent technologies, providing new insights into the nature of advanced capabilities. The model developed at the corporation level provides a more robust and consistent assessment of technology than regional-level models, particularly for finely classified technological fields. Importantly, this study shows that a corporation-level approach enables a direct comparison of technological characteristics across different regions. By focusing on corporations within specific locales, we can identify distinct industrial specializations and reveal which technologies are complex within each region’s unique context. The framework and insights from this study provide a powerful tool for creating tailored industrial policies and more accurately identifying the key technologies vital for future growth.
The rest of this paper is organized as follows. In Introduction section, we present the ubiquity and complexity of technological fields presented by the patent data, followed by a detailed analysis of technological trends. In the Finer Technological Trends and Consistency section, we provide a critical analysis of our findings and discuss the implications of our results. Finally, in the Discussion section, we detail the data processing and methods used in our analysis.
Results
Characteristics of Technologies
In this study, we use the HH algorithm to quantify the structural characteristics of technologies based on patent publications from 1,938 corporations filed between fiscal years 1981 and 2010 (see Discussion section). First, we define ubiquity (or degree centrality) $K_{T,0}$ as the number of corporations connected to a technology which lower value indicates that the technology is more rare. Similarly, the average diversity $K_{T,1}$, which measures the variety of corporations connected to a technology, implies that a higher value reflects diversified publication, whereas a lower value suggests specialization. Generally, the average of each indicator is used as a threshold to distinguish high from low values [19, 16, 20].
Another key indicator is TCI, which represents the sophistication of a technology by reflecting how advanced the knowledge required for its production. A threshold of zero divides TCI into high and low categories: a high TCI implies that a technology demands a complex combination of unique inputs and advanced technical expertise which indicates only a few advanced corporations can publish it competitively, whereas a low TCI indicates reliance on widely available capabilities, allowing many corporations to participate.
To systematically evaluate technological characteristics, we map the relationships among ubiquity $K_{T,0}$, average diversity $K_{T,1}$, and TCI across patent technologies. (Fig. 1) Technologies are classified according to the Schmoch system [21], as illustrated in Panel A of Fig. 1. Values for ubiquity and average diversity are normalized by their respective means, while TCI is segmented by the zero threshold.
We find a negligible correlation between ubiquity $K_{T,0}$ and average diversity $K_{T,1}$ (Pearson correlation coefficient $r=0.064$), indicating that widely disseminated technologies are not necessarily associated with diversified corporate portfolios (Panel B in Fig. 1). Technologies within Electrical Engineering, for example, typically show low ubiquity and low average diversity, indicating concentrated innovation within narrowly defined technological domains.
Conversely, ubiquity $K_{T,0}$ and TCI exhibit a moderate positive correlation (Pearson correlation coefficient $r=0.594$), suggesting that technologies associated with fewer corporations often require complex capabilities (Panel C in Fig.1). In this context, Electrical Engineering exhibits low ubiquity, low average diversity, and low TCI values, indicating that corporations in this field tend to concentrate their innovation efforts within narrowly defined domains.
Detailed analysis using International Patent Classification (IPC) classes further supports these observations (Fig. A1). Within the five major sectors defined by Schmoch [21], the relationship between ubiquity and average diversity is predominantly weak or moderate. Specifically, Electrical Engineering displays a strong positive correlation between ubiquity and TCI (Pearson correlation coefficient $r=0.738$), highlighting the coexistence of corporations specializing in broadly applicable technologies and those focusing on niche, specialized innovations.
In contrast to Electrical engineering technologies, Chemistry and Pharmaceuticals technologies mostly exhibit high TCI and the high average diversity (Panel D in Fig. 1, Pearson correlation coeffecience $r=0.316$). These sectors are characterized by rarity and high sophistication and are typically developed by more diversified corporations. Moreover, corporations in these domains are connected not only to fields with low ubiquity but also to those that are more common (e.g., Biotechnology, Analysis of Biological Materials, and Measurement as shown in the right quadrant of Panel C in Fig. 1). A detailed examination reveals a slight negative correlation between ubiquity and average diversity (Fig. A1), suggesting that in sectors with high technological complexity, corporations with more comprehensive technological portfolios tend to produce rare technologies. Furthermore, the IPC classes corresponding to Biotechnology and Measurement, show high TCI values in fields related to the production of sugar and yeast (C13, A21, C12), drug discovery (A61), and the preservation and management of these technologies (C07, A23). These findings are consistent with the socio-technical context of the period, particularly reflecting the emergence of drug discovery technologies in the 1980s and 1990s [22, 23]. Moreover, the significant roles of food chemistry and biotechnology during the 1990s and 2000s underscore their contribution to the development of additives that enabled safer food preservation and processing not only in Japan but globally [24, 25].
Figure 1: (A) The legend of the five classifications defined by Schmoch [21], along with an explanation of the technology codes used in (B), (C), and (D). (B) Pearson correlation between $K_{T,0}$ and $K_{T,1}$ ($r=0.064$). (C) Pearson correlation between $K_{T,0}$ and TCI ($r=0.594$). (D) Pearson correlation between TCI and $K_{T,1}$ ($r=0.316$). The black lines indicate the mean values of $K_{T,0}$ and $K_{T,1}$, and $\text{TCI}=0$ is also shown with a black line. Each dot is color-coded according to (A).
Finer Technological Trends and Consistency
The traditional regional approach measures technological complexity as a relative index reflecting regional technological capabilities, shaped by the economic advantages arising when corporations cluster within particular countries or cities, known as agglomeration economies [26]. Consequently, this method inherently captures spatial inequalities and is strongly influenced by urbanization levels in Japan. In terms of evaluating a region’s technological portfolio, this approach typically provides limited insights, either indicating binary specialization in technologies (0 or 1) or offering only the average TCI of the specialized technologies. Furthermore, finer classifications, such as IPC classes, tend to produce less consistent complexity indices compared to broader classifications like the Schmoch system [27]. Such inconsistencies appear in both regional and technological complexity evaluations due to the symmetric nature of the bipartite network structure underlying these calculations.
In this study, we propose examining technological complexity at the corporate level to overcome these limitations and achieve a more detailed understanding of regional technological portfolios. Specifically, by calculating TCI at the corporate level, we explore whether it is possible to capture more granular insights into technological capabilities within regions. By comparing TCI results derived from regional and corporate approaches, we assess how classification granularity influences the evaluation of inequality and stability of technological complexity across regions and corporations (Panel A in Fig. 2). In this study, we address these inconsistencies by comparing TCI results obtained from regional and corporate approaches, analyzing how classification granularity affects inequality and stability of complexity across regions and corporations (Panel A in Fig. 2).
The scaled TCI of the Schmoch system shows that in the regional approach 26 out of 35 classes fall in the range of 75 to 100. In the corporate approach, only 15 classes lie in that range, whereas the other classes display greater differences among technologies in the regional approach. This difference in stems from the corporate approach being able to characterize each technology in a more detailed manner. The HH algorithm defines ubiquity as the degree of a technology node and this measure depends on the number of paired nodes. The regional approach comprised roughly a few hundred region nodes, such as countries or EU cities[8, 28]. The corporate approach has corporation nodes numbering from one thousand to several ten thousand. The iterative calculation of TCI can therefore reflect a wider range of ubiquity.
The difference between the approaches also appears when using a finer classification. When the classification becomes more detailed and the number of technology nodes increases, the regional approach cannot define ubiquity sufficiently. In the case of Civil engineering the variance of TCI values for IPC classes corresponding to the same Schmoch class is larger than that for the corporate approach. The larger difference between the TCI values of the coarse classification and those of the finer classification indicates a greater inconsistency in the evaluation for the regional approach. A Wilcoxon signed rank test shows that the distribution of residuals between the TCI values of the Schmoch and IPC classes is significantly different (p value = 7.23e-14) and that the residuals for the corporate approach are smaller than those for the regional approach (Panel B in Fig. 2).
These consistent and detailed trends are observed not only over a 30 year period but also in evaluations over shorter periods such as 5 years. For example, in the IPC classes corresponding to Food chemistry, the trend of high TCI is consistent in the 2000s. Technologies with high scores as C13 and A21 and those that decline as A23 and A01 are captured in a consistent manner (SI Fig. A2).
Moreover, the corporate approach facilitates relative technological evaluations within specific regions, a capability absent in the traditional regional approach, which relies solely on the national-level presence of technologies. The corporate method identifies strategically important technologies within individual regions. For example, comparative analyses of TCI values for IPC classes among three representative prefectures, Tokyo, Osaka, and Aichi, highlight distinct local technological strengths. Aichi, notable for its concentration of major manufacturing corporations including Toyota, exhibits a technological profile markedly different from Tokyo and Osaka (Fig. 3). Technologies showing relatively higher local complexity despite lower overall national rankings underline region-specific corporate advantages.
Figure 2: (A) The TCI scores of IPC first Class (N = 124) and Schmoch (N = 35) for both prefectures and corporations. The scores have been scaled to values ranging from 0 to 100. Each IPC is plotted against the corresponding Schmoch axis, and in cases of multiple correspondences (i.e., Schmoch category “Chemical Engineering” maps to IPC classes B01–B08, C14, etc.), each pair is plotted individually. (B) The distribution of the absolute differences between TCI of IPC and of Schmoch systems.
Figure 3: Three scatter plots arranged in a single row compare the Corporate TCI in Japan (vertical axis) to the Corporate TCI in Tokyo (left panel), Osaka (middle panel), and Aichi (right panel) on the horizontal axis. Each axis is scaled from 0 to 100, and the black points represent data points for each classification or entity. The red dashed line indicates the diagonal (y = x).
Discussion
This study applies the Technological Complexity Index (TCI) to Japanese corporate patent data across various technological fields to identify sophisticated areas critical to innovation. Accurate assessment of technological complexity is essential for policymakers and corporate strategists aiming to maintain competitive advantages amidst intensifying global competition. Previous regional approaches have struggled with instability when employing detailed technological classifications, resulting in coarse and potentially misleading insights. This research provides three primary contributions. First, it demonstrates methodological robustness at the corporate level. Unlike regional approaches, corporate-level TCI retains stable technological rankings even when classifications become highly detailed. Our results reveal notable complexity in the Pharmaceuticals and Chemistry sectors, likely reflecting Japan’s industrial focus during the pharmaceutical industry’s formative period. This consistency across classification scales represents significant progress compared to previous regional analyses that lost reliability when switching from broader to finer classifications. Second, the empirical findings underscore the advantage of evaluating technological complexity at the corporate level. Prior research suggested that complexity rankings could converge due to methodological limitations of classification systems [8]. However, our corporate-level analysis maintains clear distinctions between technological fields even at granular classification levels. This evidence highlights the corporate approach’s potential to provide deeper and more accurate insights into technological dynamics within specific industries or regions. Third, our results offer practical implications for industrial policy and strategy. The identification of high-complexity technological sectors informs targeted investment and innovation strategies, especially relevant amid rapid advances in artificial intelligence and growing global competition. Corporate strategists can leverage these insights to allocate resources effectively, fostering technological leadership and competitive differentiation. Nevertheless, our study faces several limitations. Capturing dynamic shifts in technological trends requires a balance between long-term analysis, which ensures stability, and short-term assessments, which detect emerging innovations. Shortening aggregation periods introduces challenges due to the sensitivity of TCI calculations to sparse patent data, thus decreasing result stability [27]. The HH algorithm underlying TCI calculations evaluates complexity based on node connections within a bipartite network, making it vulnerable to inaccuracies when patent data are sparse. To address these challenges, future studies could integrate advanced methods such as the Fitness Complexity algorithm [13], which employs nonlinear iterative processes to better capture network structures, enhancing robustness and reliability [14, 15]. Additionally, weighting links by citation counts rather than raw patent numbers could offer a more precise representation of technological effort and impact. Future research directions include comparative robustness analyses of TCI with alternative indicators such as patent citations, application of machine learning techniques to predict technological shifts, and cross-national comparative studies. Such explorations will deepen our understanding of technological complexity and provide richer guidance for industrial development strategies. Ultimately, this research significantly advances our understanding of technological complexity at the corporate level in Japan. By clearly identifying key technology sectors, our findings inform strategic planning and policy-making, positioning Japan advantageously in an era marked by technological acceleration and intensified international competition.
Methods
Data
The dataset utilized in this study comprises registered patents submitted to the Japan Patent Office (JPO) covering fiscal years 1981 through 2010, encompassing detailed records of 3,189,536 patents filed by 64,622 corporations. Patent records adhere to the International Patent Classification (IPC) system and are further categorized according to the Schmoch classification [21].
The Schmoch system aggregates IPC codes into 35 distinct technological fields, ensuring balanced class sizes and technological consistency, and has been widely applied in previous research [29, 20, 30]. Nevertheless, the granularity of classification substantially influences analytical outcomes. Administrative classifications vary markedly in granularity and international comparability, potentially introducing analytical noise or biases if inadequately managed [8]. Employing consistent classification schemes, exemplified by the Schmoch system utilized in this study, mitigates such issues by providing balanced and coherent technological categories for robust analyses.
Traditional methods for measuring patent contributions typically employ raw patent counts [31, 20, 29, 32, 30, 17, 33]. However, such counting methods can lead to overestimation, particularly when patents involve multiple corporations or span multiple technological domains [29]. To address potential overestimations, the primary IPC class of each patent is mapped onto the corresponding Schmoch field, with patent value fractionally allocated by equally dividing it among associated corporations and technological fields. This fractional allocation prevents overestimation arising from multi-corporation ownership or cross-domain classifications.
Additionally, filtering criteria are essential to exclude extreme cases capable of distorting analytical outcomes [8, 29]. Accordingly, a filtering criterion based on patent distribution per corporation is applied in this study, retaining only the top 3% of corporations in patent output (Fig. A3). This filtering criterion excludes corporations exhibiting minimal patenting activity, thereby reducing significant analytical noise in network analyses. Following this filtering, the dataset comprises 2,485,027 patents filed by 1,938 corporations, representing approximately 93% of the original patent volume.
Utilizing these filtered data, technological complexity is based on a bipartite network connecting corporations to technological fields in which they are specialized. Although there are some arguments regarding the optimal specialization index for establishing connections within these networks [29], we follows prior research by employing the Revealed Technological Advantage (RTA) index [34]. The RTA index builds upon the revealed comparative advantage framework proposed by Balassa [35]. Specifically,The RTA index builds upon the revealed comparative advantage framework proposed by Balassa [35]. Specifically,the RTA index for a given corporation in a specific technological field measures the ratio between the share of patents held by corporation $C$ in technology $T$ and the share of patents in technology $T$ among all corporations. The numerator of the RTA index captures the fraction of patent activity devoted to technology $T$ for corporation $C$. Similarly, the denominator reflects the overall share of patent activity in technology $T$ among all corporations, computed as the sum of $W_{CT}$ over all corporations divided by the sum of $W_{CT}$ over all corporations and technologies. The RTA index is thus given by
$\mbox{RTA}_{CT}=\frac{W_{CT}}{\sum_{T}W_{CT}}\Bigg{/}\frac{\sum_{C}W_{CT}}{\sum_{CT}W_{CT}}\leavevmode\nobreak\ .$ (1)
We consider that RTA index values of 1 or greater indicate that corporation $C$ holds a prominent technological advantage in technology $T$. Based on this threshold, a binary adjacency matrix of the bipartite network $M_{CT}$ is defined by
$M_{CT}=\begin{cases}1&\text{if }\mbox{RTA}_{CT}\geq 1,\\ 0&\text{otherwise}.\end{cases}$ (2)
Technological Complexity
The binary matrix $M_{CT}$ forms the two metrics which characterize each node in the bipartite network by the sum of connections, known as degree centrality. The diversity of corporations $K_{C,0}$ quantifies the number of technologies in which a corporation is specialized, and indicated by $M_{CT}$. Similarly, the ubiquity of technologies $K_{T,0}$ represents the number of corporations that exhibit a technological advantage in a given field. These quantities are computed as
$K_{C,0}=\sum_{T}M_{CT},K_{T,0}=\sum_{C}M_{CT}\leavevmode\nobreak\ .$ (3)
Higher diversity values indicate corporations with broad technological capabilities, while lower ubiquity values signify that the technologies are rare. Through calculating these two characteristics of adjacent nodes iteratively, the diversity and ubiquity of each node is updated as
$K_{C,N}=\frac{1}{K_{C,0}}\sum_{T}M_{CT}\,K_{T,N-1},$ (4)
$K_{T,N}=\frac{1}{K_{T,0}}\sum_{C}M_{CT}\,K_{C,N-1}.$ (5)
Here, the metric $K_{T,1}$ represents the average nearest neighbor degree for technology nodes and it reflects the average diversity of corporations connected to a given technology. Substituting Eq. 4 into Eq. 5 yields
$K_{T,N}=\sum_{T^{\prime}}\widetilde{M}_{TT^{\prime}}K_{T^{\prime},N-2}$ (6)
where $T^{\prime}$ denotes a technological field, and $\widetilde{M}_{TT^{\prime}}$ is defined as
$\widetilde{M}_{TT^{\prime}}=\sum_{C}\frac{M_{CT}\,M_{CT^{\prime}}}{K_{C,0}\,K_{T,0}}\leavevmode\nobreak\ .$ (7)
Eq. 6 is satisfied when $K_{T,N}=K_{T,N-2}=1$, which corresponds to the eigenvector associated with the largest eigenvalue of the stochastic matrix $\widetilde{M}_{T,T^{\prime}}$. Since this eigenvector is a vector of ones and not informative, the eigenvector corresponding to the second largest eigenvalue of $\widetilde{M}_{TT^{\prime}}$ captures the greatest variance among technology fields [8, 36]. Thus, TCI is given by
$TCI=\frac{\widetilde{T}-\langle\widetilde{T}\rangle}{\text{stdev}(\widetilde{T})},$ (8)
where $\widetilde{T}$ represents the eigenvector corresponding to the second largest eigenvalue of $\widetilde{M}_{TT^{\prime}}$, and $\langle\widetilde{T}\rangle$ is their mean and $\text{stdev}(\widetilde{T})$ is their standard deviation.
Data Availability
The patent data utilized in this study were obtained from the Japan Patent Office (JPO) via the J-platPat portal (https://www.j-platpat.inpit.go.jp/). Records published within the most recent twelve months are directly downloadable; data older than one year can be provided by the JPO upon reasonable request. Patent classification concordance tables mapping International Patent Classification (IPC) classes to Schmoch technology fields were acquired from the World Intellectual Property Organization (WIPO) IP Statistics portal (https://www.wipo.int/en/web/ip-statistics).
Code Availability
All scripts and code used to compute the Technological Complexity Index and reproduce the results of this paper are publicly available in the following GitHub repository: https://github.com/RintaroKARASHIMA/KCIinJapaneseFirms. The implementation of the Economic Complexity algorithms is based on the py-ecomplexity library (MIT License), available at https://github.com/cid-harvard/py-ecomplexity.